OpenStudy (anonymous):

solve using the quadratic formuala. x^2-x-10=0

OpenStudy (radar):

Do you know what the quadratic looks like?

OpenStudy (radar):

*quadratic formula

OpenStudy (radar):

In case you have forgotten it. \[(-b \pm \sqrt{b ^{2}-4ac})2a\]

OpenStudy (anonymous):

yes but still i need someone to work it out for me

OpenStudy (radar):

In your equation x^2 -x-10=0 is already set up. the a, b, c in the quadratic fromula refers to the coefficients. The a is the coefficient of your variable (x^2), the b is the coefficient of the x, and c is the numerical constant. So what are the coefficients in your equation? Start with the x^2 what is its coefficient?

OpenStudy (radar):

Do some of review of using the quadratic formula. bye

OpenStudy (amistre64):

radar: nice typo :) but its..\[\frac{-b \pm \sqrt{b^2 -4ac}}{2a}\]